# Multi-objective Bayesian optimization for Likelihood-Free inference in sequential sampling models of decision making

### v2

@@ 4.2.3 Real-world data
Figure 5: Approximate and closed-form posteriors with MAPs of MLBA example. Each plot from the figure shows the marginal posterior of the parameter of interest. The green curve represents the MOBOLFI approximate posterior for joint data sources. The orange/blue curve shows the marginal MOBOLFI approximate posterior for RT /CH data, respectively. The green dot is the maximum a posteriori probability (MAP) estimate sought over all MCMC samplers of the joint MOBOLFI approximate posterior. The red curve shows the closed-form posterior conditional on both data sources, i.e. product closed-for likelihood of both data sources and a Uniform prior. The red dot denotes the MAP estimated by closed-form MLBA posterior with the same sampling strategy. The parameter $\lambda_{1}$ has been transformed to log scale to more easily compare the marginal posteriors for this parameter for different methods.
@@ 5 Discussion
Although the MOBOLFI method is motivated by the parameter inference of SSMs with multi-source data, it can be applied to Likelihood-Free inference of other complex models. For example, in Section D of the Supplement we consider an example on bacterial infection in day care centres where parameters of a latent Markov process are inferred. We partition 4 data summaries into two and apply MOBOLFI for inference. MOBOLFI is competitive with BOLFI even in this example with only an individual data source. For the case of SSMs, we use the MLBA simulator as the simulator of interest in our experiments, and MOBOLFI could be used in a wide range of SSM variants, especially for those whose likelihood functions are intractable like DDM, MDFT, etc. An interesting direction for future work is to consider different choices of the acquisition function in the MOBOLFI approach. Multi-objective BO is an active area of research in Bayesian optimization, and new developments in BO methodology may translate into improved performance in Likelihood-Free applications. State-of-art computational methods for the MLBA with closed-form likelihood, including variants of the model including random effects, are discussed in Gunawan et al. (2020), and Evans (2019) addresses the important issue of efficient simulation from various SSM models, which is particularly important for Likelihood-Free inference.
In applying MOBOLFI there are a number of challenges. One is that the initialization of the algorithm can be very important, and an adequate initial covering of the space is needed. This seems to be especially crucial with acquisition functions based on expected hypervolume improvement, where the BO algorithm can keep proposing points in a small area of the space without a proper initialization. Another difficulty, common to many other LFI algorithms, is the choice of data summary statistics used to define discrepancies. In Section C.4 of the Supplement we consider alternative choices of summary staistics to the ones considered in section 4.2.2 based on auxiliary model approaches to summary statistic construction which are widely used in the ABC literature (e.g. Drovandi, Pettitt and Lee 2015). Further exploration of these challenges is left to future work.
@@ A.3 Implementation of MOBOLFI
One classic scaling method in the machine learning literature is normalization by using Mean absolute deviation (MAD). Algorithm 1 presents the detailed steps of doing scaling of a joint noisy objective $\Delta$ for implementing BOLFI. They key idea of Algorithm 1 is to put each elements of $\Delta$ on a similar scale. The use of Algorithm 1 does not always result in the best performance. In the plot 6(a), the green curve with $w=0.7$ os chosen by Algorithm 1 (rounding to 1 decimal places). The scaling with best performance is example specific, depending on the information of each data source, the LFI method and the BO acquisition function. We leave further investigation of optimal scaling approaches to future work. For the results of this paper, by default we choose the scaling as the outcome (rounded to 1 decimal places) from algorithm 1 using $n=100$. The auxiliary model example (orange curve in Figure 11) defines a discrepancy as the score vector of an auxiliary model, and the scaling in this example is slightly different and will be discussed in Section C.3 of this supplement.
@@ B.2 Extra findings - toy example
We apply MOBOLFI with the same experiment setup as in Section B.1 of this supplement. The MOBOLFI and BOLFI approximate posteriors are presented in Figure 8. We focus on the parameters $\theta_{9},\theta_{10}$, which are parameters influencing the distribution of only one of the data soruces. In plot 8(a), the MOBOLFI approximate posterior, leveraging correlated noise between data sources, still outperforms BOLFI in inference of $\theta_{9},\theta_{10}$. In plot 8(b), for $\theta_{10}$ not depending on $X$, the MOBOLFI approximate posterior conditional on of $X$ does not obtain smaller variance than the BOLFI approximate posterior. That is not surprising to see, since information from $X$ is viewed as useless and redundant for inference of $\theta_{10}$. On the other hand, MOBOLFI still performs better than BOLFI in the inference of $\theta_{9}$. Plot 8(c) shows that approximate MOBOLFI posterior conditional on $W$ only does not obtain smaller variance than the BOLFI approximate posterior, for $\theta_{9}$ not depending on $W$.
@@
Figure 8: Approximate posterior for the updated toy example with two parameters $\theta_{9},\theta_{10}$ solely depending on two data sources $X,W$, respectively. The row 8(a) shows the MOBOLFI and BOLFI approximate posteriors under the updated simulator, colored by blue/orange. The green curve is the real posterior, the red curve is the prior and the dash red line is the true value of $\theta_{9},\theta_{10}$. In row 8(b)/8(c) the MOBOLFI and BOLFI approximate posteriors calculated by marginal likelihood of $X/W$ are given by blue/orange lines. The green curve is the real posterior of $X/W$ respectively.
@@ C.4 Extra findings - MLBA example (empirical case)
C.4.1 MLBA variants selection
In the main text,
$\theta=(\lambda_{1},\delta_{ICEV},\log{\beta_{TC}},\log{\beta_{\log{DR}}},\log{(\chi-\mathcal{A})})$
is estimated fixing $\lambda_{2}=0$. Theoretically, $\lambda_{2}$ is also estimable. Table 1 here provides more justification for selecting the $\lambda_{2}=0$ MLBA variant for further study on MOBOLFI. For MLBA variant with free $\lambda_{2}$, its MAP value is very close to 0. Fixing $\lambda_{2}=0$ not only reduces the number of estimable parameters but brings a slightly higher log-likelihood.
Table 1: Closed-form parameter estimate based on two MLBA variants.
$\lambda_{2}=0$ $\lambda_{2}$ is free
parameter
posterior mean
posterior MAP
(95% HDI)
posterior mean
posterior MAP
(95% HDI)
$\log{\beta_{TC}}$
2.055
2.082
( 1.848 , 2.264 )
2.179
2.068
( 1.911 , 2.462 )
$\log{\beta_{\log{DR}}}$
1.141
1.128
( 0.814 , 1.49 )
1.438
1.116
( 0.893 , 2.701 )
$\delta_{ICEV}$
0.431
0.55
( -0.369 , 1.145 )
-0.005
0.511
( -1.712 , 1.114 )
$\lambda_{1}$
0.199
0.203
( 0.115 , 0.306 )
0.196
0.202
( 0.132 , 0.29 )
$\lambda_{2}$
0.073
0.001
( 0 , 0.167 )
$\log{(\chi-\mathcal{A})}$
0.705
0.752
( 0.463 , 0.921 )
0.689
0.73
( 0.452 , 0.918 )
Log-likelihood -2507.614 -2507.52 -2509.831 -2509.107
C.4.2 Results with other discrepancy design
In the main text, we used an average of 1-Wasserstein distance over S = 30 simulated datasets for RT discrepancy $\Delta^{*}_{1}$ and Brier Score for CH discrepancy $\Delta^{*}_{2}$ for the empirical data experiment. As stated there, the chosen discrepancies are more stable and informative compared to those based on a single model simulation, $\Delta_{1}$ and $\Delta_{2}$. Figure 14 shows the MOBOLFI posterior comparison for two different discrepancy designs. According to De-MCMC sampling of the posterior, $\Delta_{1}$ and $\Delta_{2}$ (eq.4.4 and 4.5 in the main text) bring larger uncertainty to the posterior compared to $\Delta^{*}$. Admittedly, there is a trade-off between discrepancy computation speed ($\Delta^{*}$ takes S times as much time as $\Delta$ does for the same number of BO acquisitions) and the number of BO acquisitions required for good likelihood approximation. Overall, $\Delta^{*}$ offers a quicker convergence for MOBOLFI training and smaller uncertainty in its approximate posteriors.
Figure 14: Approximate posteriors of MOBOLFI with different discrepancy designs. The dashed line stands for the MOBOLFI with discrepancy design $\Delta$ the same as the synthetic data example. Green is the joint approximate posterior for both CH and RT data, while orange and blue stands for the marginal posterior by RT/CH data only.
C.4.3 Weak dependence of discrepancies on $\beta$ and $\lambda_{1}$
Figure 15 shows histograms of $1000$ discrepancies for simulation data generated by both MOBOLFI and closed-form MAPs. Data generated by two MAPs results in similar distributions of discrepancies, despite closed-form MAP has values of $\log{\beta_{TC}},\log{\beta_{\log{DR}}}$ and $\lambda_{1}$ far from MOBOLFI MAP. This indicates that with appropriately selected $\delta_{ICEV}$ and $\log{(\chi-\mathcal{A})}$, the discrepancies (and the corresponding MLBA simulation) in this experiment are insensitive to $\log{\beta_{TC}},\log{\beta_{\log{DR}}}$ and $\lambda_{1}$.
Figure 15: Discrepancy distributions for MAPs over 1000 repetitions (MOBOLFI MAP is in green and closed-form MAP is in red, respectively). The MAPs are selected with the largest (approximated) posterior value among all De-MCMC samplers for posterior estimation.
Figure 16: Boxplots for discrepancies over the domain of each parameter while others are fixed to MAP value. The boxplot with the orange line shows the discrepancies at MAP point.
To further investigate the effect of $\log{\beta_{TC}},\log{\beta_{\log{DR}}}$ and $\lambda_{1}$, a sensitivity analysis is done to evaluate changes on discrepancies depending on change on each parameter. Figure 16 demonstrates the discrepancy values over each parameter domain when others are fixed to the MOBOLFI MAP. For each parameter 14 points are selected, of which one is the MAP value and the other 13 are boundaries and its 0.1 neighbour points for 5 equally divided segments over the domain. Take $\delta_{ICEV}$ as an example, $\{-3,-2.9,-1.1,-1,-0.9,0.56,0.9,1,1.1,2.9,3,3.1,4.9,5\}$ are selected while 0.56 is the MAP value, $\{-3,-1,1,3,5\}$ are the boundaries of five equal segments over domain $[-3,5]$ and the rest are associated $0.1$ neighbor points within the domain. For each parameter, $100$ independent replicates of discrepancy calculation are presented in boxplots. Compared to $\delta_{ICEV}$ and $\log\{(\chi-\mathcal{A})\}$, $\log{\beta_{TC}},\log{\beta_{\log{DR}}}$ and $\lambda_{1}$ have a weak impact on discrepancies when values of them vary dramatically over the domain. This claims that with appropriately selected $\delta_{ICEV}$ and $\log{(\chi-\mathcal{A})}$, the discrepancies (and the corresponding MLBA simulation) in this experiment weakly depend on $\log{\beta_{TC}},\log{\beta_{\log{DR}}}$ and $\lambda_{1}$.

### v3

@@ 4.2.3 Real-world data
Figure 5: Approximate and closed-form posteriors with posterior mean for MLBA example. Each plot from the figure shows the marginal posterior of the parameter of interest. The green curve represents the MOBOLFI approximate posterior for joint data sources. The orange/blue curve shows the marginal MOBOLFI approximate posterior conditional on RT /CH data, respectively. The green dot is the posterior mean. The red curve shows the closed-form posterior conditional on both data sources, i.e. product closed-form likelihood of both data sources and a Uniform prior. The red dot denotes the posterior mean of closed-form MLBA estimated with the same sampling strategy.
Figure 5 shows approximate marginal MOBOLFI posteriors for selected parameters and the corresponding posteriors for the closed-form MLBA likelihood. The figure demonstrates that the contributions of different data sources to the estimation of individual parameters can vary greatly according to the parameter, and MOBOLFI can reveal these relationships. On the one hand, the $CH$ data contributed more to the estimation of the front-end parameters of MLBA ($\delta_{ICEV},\mathbf{\beta},\lambda_{1}$) related to the preference across alternatives and attributes, while the $RT$ data contributes more to estimation of the back-end parameters of MLBA ($\log{(\chi-\mathcal{A})}$). For example, the MOBOLFI marginal posterior of $\lambda_{1}$ and $\log{(\beta_{DR})}$ is very diffuse when conditioning on only response-time data, showing that the MOBOLFI posterior conditioning on joint data sources is informative mostly because of the choice data. In contrast, the $\log{(\chi-\mathcal{A})}$ MOBOLFI marginal posterior conditioning on both choice and response time is more similar to its marginal posterior conditional on response time only. This is consistent with the interpretation that $\chi$ reflects the decision difficulty and respondent’s cautiousness, which are associated with RT. On the other hand, the approximate MOBOLFI posterior conditional on both data sources is always less dispersed than the MOBOLFI posterior conditional on a single data source only. This observation is aligned with the finding (Li and Bansal, 2024) that joint choice-RT data provides a more efficient parameter estimate compared to an estimate applying marginal data sources (CH or RT only).
The posterior estimates of MOBOLFI are different from those of closed-form in Figure 5, mainly on $\log{\beta_{TC}}$ and $\delta_{ICEV}$. Figure 6 examines how well the observed data are reproduced in the fitted model for these point estimates. The left column in the figure shows that log discrepancies for synthetic data simulated using the posterior mean of MOBOLFI are generally larger than those for the closed-form likelihood posterior mean value on both $CH$ and $RT$. The right column of the figure compares simulated discrepancies for MOBOLFI and close-form MAP point estimates. The $RT$ log discrepancies simulated using MOBOLFI MAP are smaller, while $CH$ log discrepancies simulated using closed-form MAP are smaller.
Figure 6: The histogram of discrepancies with 1000 replicates based on different point estimates. The MAP of MOBOLFI has slight distance from its post mean mainly on $\delta_{ICEV}$ and $\log{(\beta_{TC})}$, whose value can be checked in Section C.4 of the Supplement.
@@ 5 Discussion
Although the MOBOLFI method is motivated by the parameter inference of SSMs with multi-source data, it can be applied to Likelihood-Free inference of other complex models. For example, in Section D of the Supplement we consider an example on bacterial infection in day care centres where parameters of a latent Markov process are inferred. We partition 4 data summaries into two and apply MOBOLFI for inference. MOBOLFI is competitive with BOLFI even in this example with only an individual data source. For the case of SSMs, we use the MLBA simulator as the simulator of interest in our experiments, and MOBOLFI could be used in a wide range of SSM variants, especially for those whose likelihood functions are intractable like DDM, MDFT, etc. An interesting direction for future work is to consider different choices of the acquisition function in the MOBOLFI approach. Multi-objective BO is an active area of research in Bayesian optimization, and new developments in BO methodology may translate into improved performance in Likelihood-Free applications. State-of-the-art computational methods for the MLBA with closed-form likelihood, including variants of the model including random effects, are discussed in Gunawan et al. (2020). Evans (2019) addresses the important issue of efficient simulation from various SSM models, which is particularly important for likelihood-free inference.
In applying MOBOLFI there are a number of challenges. One is that the initialization of the algorithm can be very important, and an adequate initial covering of the space is needed. This seems to be especially crucial with acquisition functions based on expected hypervolume improvement, where the BO algorithm can keep proposing points in a small area of the space without a proper initialization. Another difficulty, common to many other LFI algorithms, is the choice of data summary statistics used to define discrepancies. In Section C.3 of the Supplement we consider alternative choices of summary statistics to the ones considered in section 4.2.2 based on auxiliary model approaches to summary statistic construction which are widely used in the ABC literature (e.g. Drovandi, Pettitt and Lee 2015). The choice of informative summary statistics remains challenging when working with complex models like SSMs. Further exploration of these issues is left to future work.
@@ A.3 Implementation of MOBOLFI
One classic scaling method in the machine learning literature is normalization by using Mean absolute deviation (MAD). Algorithm 1 presents the detailed steps of doing scaling of a joint noisy objective $\Delta$ for implementing BOLFI. They key idea of Algorithm 1 is to put each elements of $\Delta$ on a similar scale. The use of Algorithm 1 does not always result in the best performance. In the plot 7(a), the green curve with $w=0.7$ os chosen by Algorithm 1 (rounding to 1 decimal places). The scaling with best performance is example specific, depending on the information of each data source, the LFI method and the BO acquisition function. We leave further investigation of optimal scaling approaches to future work. For the results of this paper, by default we choose the scaling as the outcome (rounded to 1 decimal places) from algorithm 1 using $n=100$. The auxiliary model example (orange curve in Figure 12) defines a discrepancy as the score vector of an auxiliary model, and the scaling in this example is slightly different and will be discussed in Section C.3 of this supplement.
@@ B.2 Extra findings - toy example
We apply MOBOLFI with the same experiment setup as in Section B.1 of this supplement. The MOBOLFI and BOLFI approximate posteriors are presented in Figure 9. We focus on the parameters $\theta_{9},\theta_{10}$, which are parameters influencing the distribution of only one of the data soruces. In plot 9(a), the MOBOLFI approximate posterior, leveraging correlated noise between data sources, still outperforms BOLFI in inference of $\theta_{9},\theta_{10}$. In plot 9(b), for $\theta_{10}$ not depending on $X$, the MOBOLFI approximate posterior conditional on of $X$ does not obtain smaller variance than the BOLFI approximate posterior. That is not surprising to see, since information from $X$ is viewed as useless and redundant for inference of $\theta_{10}$. On the other hand, MOBOLFI still performs better than BOLFI in the inference of $\theta_{9}$. Plot 9(c) shows that approximate MOBOLFI posterior conditional on $W$ only does not obtain smaller variance than the BOLFI approximate posterior, for $\theta_{9}$ not depending on $W$.
@@
Figure 9: Approximate posterior for the updated toy example with two parameters $\theta_{9},\theta_{10}$ solely depending on two data sources $X,W$, respectively. The row 9(a) shows the MOBOLFI and BOLFI approximate posteriors under the updated simulator, colored by blue/orange. The green curve is the real posterior, the red curve is the prior and the dash red line is the true value of $\theta_{9},\theta_{10}$. In row 9(b)/9(c) the MOBOLFI and BOLFI approximate posteriors calculated by marginal likelihood of $X/W$ are given by blue/orange lines. The green curve is the real posterior of $X/W$ respectively.
@@ C.4 Extra findings - MLBA example (empirical case)
Figure 15 shows the difference between the posterior mean and MAP estimates from the MOBOLFI approximate posterior and closed-form true posterior. The largest differences are for $\delta_{ICEV}$ and $\log{\beta_{TC}}$. Point estimates from the MOBOLFI approximate posterior have smaller value of $\hat{\delta}_{ICEV}$ and larger value on $\log{\hat{\beta}_{TC}}$, when compared to estimates from closed-form posterior.
Figure 15: The posterior density and point estimates of MOBOLFI and closed-form based on CH and RT data sources.
Appendix D Day care center example